OpenStudy (ayaan):

-2x^2-20x+16 middle term break please help

OpenStudy (gorv):

sum of middle term = -20 and product of trem broken = -32

OpenStudy (gorv):

can u do that??

OpenStudy (ayaan):

can you please solve?

OpenStudy (gorv):

i can guide

OpenStudy (gorv):

well use quadratic formula

OpenStudy (gorv):

u know abt that??

hartnn (hartnn):

are you sure about the question ? you won't be getting integer solutions so you cannot split middle term

OpenStudy (ayaan):

yeah i know about it

OpenStudy (gorv):

can u apply it for finding x

hartnn (hartnn):

use quadratic formula or completing the square method whichever you're comfortable with

OpenStudy (ayaan):

but i have forgotten the quadratic formula

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (ayaan):

can you guide me how to put values in that

hartnn (hartnn):

first find a,b,c to simplify calculations, you can also factor -2 initially from your equation

OpenStudy (ayaan):

okay i am working on it

OpenStudy (ayaan):

thnx for your help

hartnn (hartnn):

\(-2 (x^2+10x-8)\) compare x^2+10x-8 with ax^2 +bx+c

OpenStudy (ayaan):

a=-2, b=20,c=8

hartnn (hartnn):

no.. you're taking x^2+10x-8 right ?

OpenStudy (ayaan):

yeah

hartnn (hartnn):